A rod of mass M kg and length L m is bent in the form of an equilateral triangle as shown. The MI of traingle about a vertical axis perpendicular to the plane of triangle and passing through the centre is?

Grade:12

6 Answers

Vikas TU
10 years ago

 

Let a be the side of the equilateral triangle.

Then, L = 3a or a = L/3

        thus, M is the mass of the rod the then the three sides of the triangle will be of ''M/3 mass.

and also, by construction, O is the centre or Centroid of the triangle. and Centroid is the points of intersectin of all the medians of the triangle. therefore,

 

tan30'' = 1/root3 = 6p/l or p = l/6root3 where p is the perpendicular drawn on side AB from point O.

Moment of Inertia of traingle about a vertical axis perpendicular to the plane of triangle and passing through the centre is :

I = [M/3 x L/6root3]^2 x 3 (other three rods of same dimension.]

I = Ml^2/108

is the ansr.

 

Plz Approve!



Roshan Mohanty
10 years ago

Is it M.L^2 / 2???

Shuvam Shukla
10 years ago

ans is ML^2/54..

Roshan Mohanty
10 years ago

Sorry some calc mistake 
i am getting ML^2 / 18 

Vikas TU
10 years ago

I dont know then I checked twice but getting the same anr.

vishal ojha
10 years ago

ml^2/108...